OpenStudy (anonymous):

The value, V, of a Tiffany lamp, worth $225 in 1965, increases at 15% per year. Its value in dollars t years after 1965 is given by V=225(1.15)^t. Find the average value of the lamp over the period 1965-2000.

OpenStudy (anonymous):

@Dockworke I'm pretty horrible at typing and explaining, so I'm sort of relying to people back me up :S

OpenStudy (anonymous):

i think you should subtract 225 from v(35) before dividing by 35

OpenStudy (anonymous):

Yeah, delta y/ delta x, make sense!

OpenStudy (anonymous):

this may be an integration problem though

OpenStudy (anonymous):

Avg = Δ V / Δ t = V ( 35 ) - 35 / 35 = [ 225(1.15)^ 35 - 35 ] / 35 = $849.7

OpenStudy (anonymous):

I figured it out on my own. The answer is actually $6079.62, but thank you for having tried.